# Quantum sampling on a quantum annealer for large volumes in the strong coupling limit for gauge group U(3)

### v1

@@ II Effective theory of Lattice QCD
Our effective theory is derived from staggered fermions in the strong coupling limit, which is well known to have a dual representation in terms of integer-valued dual variables [13]. These dual variables consist of so-called monomers $m(n)$ and dimers $d_{\nu}(n)$ (with $n$ a lattice coordinate) for ${\rm U}(3)$ gauge theory, and also baryonic world-lines, which have to be included for ${\rm SU}(3)$ gauge theory. We have studied $SU(3)$ in [2] but in this work we restrict ourselves to ${\rm U}(3)$ for simplicity. As explained detail in [1], the integer-valued dual variables can be mapped to a binary vector $x$ which can be adapted for the quadratic unconstrained binary optimization (QUBO) formalism:
@@ III Technical development
Figure 1: Selecting $2\times 2$ sub volume with fixed boundary condition. Black dots are monomers, blue lines are dimers. Monomers and dimers inside of red boxes are updated in parallel. The Grassmann constraint is required be satisfied at each site.
@@ III.3 Metropolis and Metropolis-Hastings
Figure 6: Comparison of the acceptance rate Left: for a $4\times 4$ lattice from Metropolis-Hastings, with $2\times 2$ sub-lattice histograms from D-Wave, for various penalty factors $p$. Center: illustrating independence of the acceptance rate of the volume. Right: Metropolis from classical computation, for $\gamma=0.1,1.0$. Metropolis-Hastings for $p=\infty$ reproduces the classical Metropolis acceptance rate.
@@ III.4 Branching Strategy
Unfortunately a limitation of this method so far is that the memory requirement to branches grows exponentially with $N_{\rm branch}$, preventing us from presenting results at this point in time.
@@ V Conclusion
Extending to ${\rm SU}(3)$ and including gauge corrections are another important next steps towards a more realistic effective theory of lattice QCD. The first aspect includes baryons, the second aspect includes gluon propagations. The dual representation is well established for both extensions, and again for can be mapped on binary vectors. However, these formulations will require more logical qubits. Still, it remains feasible to map the logical qubits required for the QUBO matrix to the physical qubits on the annealer. However, while the histograms required for this work, given square sampling, had $4^{4}$ different boundaries to classify the histograms, the generalizations require up to $16^{4}$ different boundaries, for which the histograms can no longer be pre-computed on D-wave. Instead they need to be computed during Metroplis-Hastings in between the updates, whenever required. Studying more realistic effective theory will introduce two more physical parameters [14]: the quark chemical potential $\mu_{q}$, and the inverse gauge coupling $\beta$ that is related to the lattice spacing $a$. We will report on our findings using the same approach via histograms determined on the annealer in a subsequent publication.

### v2

@@ II Effective theory of Lattice QCD
Our effective theory is derived from staggered fermions in the strong coupling limit, which is well known to have a dual representation in terms of integer-valued dual variables [13]. These dual variables consist of so-called monomers $m(n)$ and dimers $d_{\nu}(n)$ (with $n$ a lattice coordinate) for ${\rm U}(3)$ gauge theory, and also baryonic world-lines, which have to be included for ${\rm SU}(3)$ gauge theory. We have studied ${\rm SU}(3)$ in [2] but in this work we restrict ourselves to ${\rm U}(3)$ for simplicity. As explained detail in [1], the integer-valued dual variables can be mapped to a binary vector $x$ which can be adapted for the quadratic unconstrained binary optimization (QUBO) formalism:
@@ III Technical development
Figure 1: Selecting $2\times 2$ sub volume with fixed boundary condition. Black dots are monomers, blue lines are dimers. Monomers and dimers inside of red boxes are updated in parallel. The Grassmann constraint is required to be satisfied at each site.
@@ III.3 Metropolis and Metropolis-Hastings
Figure 6: Comparison of the acceptance rate. Left: for a $4\times 4$ lattice from Metropolis-Hastings, with $2\times 2$ sub-lattice histograms from D-Wave, for various penalty factors $p$. Center: illustrating independence of the acceptance rate of the volume. Right: Metropolis from classical computation, for $\gamma=0.1,1.0$. Metropolis-Hastings for $p=\infty$ reproduces the classical Metropolis acceptance rate.
@@ III.4 Branching Strategy
Unfortunately a limitation of this method so far is that the memory requirement for branching grows exponentially with $N_{\rm branch}$, preventing us from presenting results at this point in time.
@@ V Conclusion
Extending to ${\rm SU}(3)$ and including gauge corrections are another important next step towards a more realistic effective theory of lattice QCD. The first aspect includes baryons while the second aspect includes gluon propagation. The dual representation is well established for both extensions, and again can be mapped on binary vectors. However, these formulations will require more logical qubits. Still, it remains feasible to map the logical qubits required for the QUBO matrix to the physical qubits on the annealer. However, while the histograms required for this work, given square sampling, had $4^{4}$ different boundaries to classify the histograms, these generalizations will require up to $16^{4}$ different boundaries, for which the histograms can no longer be pre-computed on D-wave. Instead they need to be computed during Metroplis-Hastings in between the updates, whenever required. Studying more realistic effective theories will introduce two more physical parameters [14]: the quark chemical potential $\mu_{q}$, and the inverse gauge coupling $\beta$ that is related to the lattice spacing $a$. We will report on our findings using the same approach via histograms determined on the annealer in a subsequent publication.